Operation Summary for May 30, 2011 - June 1, 2011 |
|
Monday May 30 The day shift began with Operations monitoring the Tevatron (TEV) store 8781 (274.65E30) and the Recycler stash (154.6E10), with the Antiproton Source (Pbar) stacking (8.2E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off for target replacement, with the MiniBooNE and Meson FTBF T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, operators reported that the MI-52 kicker tripped off at 2 PM and would not reset. Operators called in kicker personnel to investigate. The Booster low level RF (LLRF) room temperature got too warm, which caused problems with the Booster RF, and cut off beam to MiniBooNE. FESS personnel investigated the temperature problem and found an air conditioner off; they restarted it. Operators had beam restored to MiniBooNE by the end of shift. During the evening shift, Operations disabled beam to T-978 at 6:14 PM. MI kicker personnel made adjustments to the MI-52 kicker and brought it back online. Operations terminated store 8781 at 10:18 PM, put the TEV into HEP shot setup, and established store 8782 at 11:58 PM with an Average Initial Luminosity (AIL) of 373.16E30. Operators explained that shot setup took a little longer due to some issues with Electron Cooling. Tuesday May 31 The midnight shift began with Operations monitoring store 8782 (332.68E30) and the stash (60.2E10), with Pbar stacking (9.7E10), and with MiniBooNE taking beam. During the midnight shift, Operations established beam to T-978 at 4:30 AM. Operators reported beam position problems with the MI-52 extraction on the Switchyard event; operators tuned to resolve these problems. Operators reported some short downtime due an MI-8 corrector magnet bulk power supply tripping off. During the day shift, Operations established beam to MTA at 2:15 PM. At 2:39 PM, a lightning strike caused the loss of store 8782 due to a fast bypass interlock loop drop out. There was no quench. Operators reported finding a bad power supply at TEV sector D4; EE Support personnel replaced the supply. TEV recovery continued through the end of the shift. A vacuum burst in the MI RF line caused about 20 minutes of downtime. A sequencer ACL script error caused a 22E10 loss of antiprotons during a transfer. Problems with Linac and Booster RF systems caused other short down times. During the evening shift, MTA took beam until 4:37 PM. Operations disabled beam to T-978 at 5:27 PM. Operations had the TEV ready for shot setup by 5:30 PM and established store 8784 at 7:31 PM with an AIL of 397.77E30. Kicker personnel worked on the MI-52 kicker, which caused a few minutes of downtime. Wednesday June 1 The midnight shift began with Operations monitoring store 8784 (168.07E30) and the stash (163.7E10), with Pbar stacking (15.7E10), and with MiniBooNE taking beam. During the midnight shift, Operations established beam to T-978 at 4:17 AM. Operators halted Linac beam at 7:15 AM to prepare Linac and Booster for access. Operations terminated store 8784 at 7:35 AM. Operators prepared CDF, D0, the TEV Transfer Hall, and enclosures A-E for accesses. Operations reported short down times caused by Booster and Linac RF system trips. The Plans for Wednesday The plans for today are to access the TEV for cryo system wet engine repairs and other maintenance work. All TEV work should be completed by late afternoon. The Booster and Linac work should be completed by late morning. The Run Coordinator expects the next shot setup to begin on the evening shift. |
|
More Information Comments and Suggestions |
| last modified 06/01/2011 email Fermilab |
| Security, Privacy, Legal | |